# SmartEraser: Remove Anything from Images using Masked-Region Guidance

### v1

@@ Abstract
Object removal has so far been dominated by the “mask-and-inpaint” paradigm, where the masked region is excluded from the input, leaving models relying on unmasked areas to inpaint the missing region. However, this approach lacks contextual information for the masked area, often resulting in unstable performance. In this work, we introduce SmartEraser, built with a new “removing” paradigm called Masked-Region Guidance. This paradigm retains the masked region in the input, using it as guidance for the removal process. It offers several distinct advantages: (a) it guides the model to accurately identify the object to be removed, preventing its regeneration in the output; (b) since the user mask often extends beyond the object itself, it aids in preserving the surrounding context in the final result. Leveraging this new paradigm, we present Syn4Removal, a large-scale object removal dataset, where instance segmentation data is used to copy and paste objects onto images as removal targets, with the original images serving as ground truths. Experimental results demonstrate that SmartEraser significantly outperforms existing methods, achieving superior performance in object removal, especially in complex scenes with intricate compositions.
@@ 1 Introduction
With the rapid advancement of generative models [35, 10, 40, 11, 15, 37, 34, 32], image editing has garnered significant attention for its broad applications. Among various editing tasks, object removal [55, 9] is especially prominent, empowering users to seamlessly eliminate unwanted elements while maintaining the realism of the original image. This functionality, essential for removing distracting people or objects in photos, has become a key feature in widely used applications like Photoshop, Google Photos, Microsoft Designer, etc.
Equal contribution.  $\dagger$ Corresponding authors.
@@
Figure 2: Comparison of the “mask-and-inpaint” paradigm and the proposed Masked-Region Guidance paradigm for object removal. Unlike the mask-and-inpaint approach, our paradigm uses the masked regions as critical guidance for object removal.
@@
Using the Masked-Region Guidance paradigm and the Syn4Removal dataset, we design a framework based on text-to-image diffusion models for object removal. To enhance the model’s robustness to varying mask shapes from user input, we introduce a mask enhancement technique that simulates different mask shapes over removal targets. Additionally, we incorporate CLIP-based visual guidance to assist the model in better understanding the removal targets. The resulting model called SmartEraser outperforms previous methods significantly in both quantitative and qualitative evaluations. We choose the name SmartEraser because the model can “smartly” identify removal targets and remove it while preserving its surrounding region.
@@ 3 SmartEraser
The goal of object removal is to completely erase the target object specified by a masked region and seamlessly generate consistent content based on the surrounding context. Most existing methods follow a “mask-and-inpaint” paradigm, where the masked area is excluded from the input, and the model attempts to inpaint the missing content based on the unmasked regions. While these methods can successfully remove objects in simple scenarios, they often regenerate new objects or produce blurs and artifacts in the masked region in more complex scenes.
In contrast, we propose a novel paradigm: Masked-Region Guidance. Instead of discarding the masked region, we retain it in the input, leveraging it to guide the removal process. This approach is based on the insight that, by knowing the exact removal targets, the model can precisely erase them while leaving the surrounding context maintained. In the following sections, we will first introduce the method of Masked-Region Guidance in Sec. 3.1. Then, we present Syn4Removal, a large-scale synthetic dataset designed for this paradigm in Sec. 3.2. Finally, we describe SmartEraser, a fine-tuned diffusion model based on the Mask-Region Guidance paradigm and Syn4Removal dataset in Sec. 3.3.
@@ 3.1 Masked-Region Guidance
Given an input image $\bm{x}$, and user-specified mask $\bm{m}$ containing the removal targets, the object removal task aims to produce the removal result $\bm{\hat{x}}$ while maintaining the realism of the original image. The mask $\bm{m}$ is a binary mask where 1 indicates the region containing the removal targets. Existing models [39, 35, 9, 55] are mostly based on the “mask-and-inpaint” paradigm, i.e., utilizing the unmasked region $\bm{x}\odot(\bm{1}-\bm{m})$ and mask $\bm{m}$ to predict the results without removal object in the masked region. These models are learned to maximize the following objective:
@@ 4.2 Comparasion with Previous Methods
ZITS++ [4] 29.257 0.838 0.820 0.388 0.522 15.442
MAT [24] 78.809 1.494 0.628 0.433 0.495 13.322
LaMa [39] 20.034 0.216 0.916 0.348 0.557 16.383
BLD [1] 65.003 1.760 0.834 0.471 0.467 15.775
RePaint [30] 51.709 1.639 0.812 0.381 0.510 15.861
SD-Inpaint [35] 38.675 0.324 0.857 0.369 0.537 16.111
CLIPAway [9] 20.253 0.123 0.915 0.333 0.577 17.434
PowerPaint [55] 21.227 0.294 0.926 0.308 0.602 18.100
SmartEraser 10.901 0.092 0.937 0.276 0.612 19.994
@@
ZITS++ [4] 5.734 0.229 0.899 0.350 0.634 19.453
MAT [24] 6.651 0.321 0.742 0.407 0.575 14.541
LaMa [39] 3.935 0.201 0.926 0.351 0.683 21.735
BLD [1] 10.755 0.746 0.913 0.418 0.632 19.346
RePaint [30] 5.145 0.822 0.924 0.325 0.644 20.607
SD-Inpaint [35] 5.651 0.196 0.912 0.320 0.656 21.691
CLIPAway [9] 4.518 0.193 0.925 0.323 0.666 21.953
PowerPaint [55] 3.769 0.275 0.934 0.278 0.715 23.870
SmartEraser 3.496 0.106 0.949 0.257 0.734 25.363
@@
ZITS++ [4] 4.007 0.193 0.810 0.358 0.581 17.315
MAT [24] 9.276 0.300 0.700 0.374 0.570 13.357
LaMa [39] 4.039 0.097 0.932 0.325 0.606 19.319
BLD [1] 9.035 0.705 0.896 0.464 0.525 16.539
RePaint [30] 5.560 0.696 0.908 0.378 0.553 17.689
SD-Inpaint [35] 5.905 0.107 0.897 0.348 0.572 18.105
CLIPAway [9] 3.842 0.093 0.930 0.344 0.593 18.561
PowerPaint [55] 4.107 0.200 0.931 0.291 0.663 20.594
SmartEraser 2.634 0.053 0.939 0.269 0.672 22.029
@@ 4.3 Ablation studies
Effect of the Masked-Region Guidance paradigm: We observe that the baseline model often regenerates objects instead of properly removing them, as it is trained in a traditional “inpainting” manner. Incorporating masked-region guidance significantly improves the removal performance by reducing the risk of object regeneration. However, due to the mismatch between the mask shape used during training and the user-defined mask during testing, the model does not always perfectly remove the target objects and preserve the surrounding context.

### v2

@@ Abstract
Object removal has so far been dominated by the “mask-and-inpaint" paradigm, where the masked region is excluded from the input, leaving models relying on unmasked areas to inpaint the missing region. However, this approach lacks contextual information for the masked area, often resulting in unstable performance. In this work, we introduce SmartEraser, built with a new “removing" paradigm called Masked-Region Guidance. This paradigm retains the masked region in the input, using it as guidance for the removal process. It offers several distinct advantages: (a) it guides the model to accurately identify the object to be removed, preventing its regeneration in the output; (b) since the user mask often extends beyond the object itself, it aids in preserving the surrounding context in the final result. Leveraging this new paradigm, we present Syn4Removal, a large-scale object removal dataset, where instance segmentation data is used to copy and paste objects onto images as removal targets, with the original images serving as ground truths. Experimental results demonstrate that SmartEraser significantly outperforms existing methods, achieving superior performance in object removal, especially in complex scenes with intricate compositions.
@@ 1 Introduction
With the rapid advancement of generative models [35, 10, 40, 11, 15, 37, 34, 32], image editing has garnered significant attention for its broad applications. Among various editing tasks, object removal [55, 9] is especially prominent, empowering users to seamlessly eliminate unwanted elements while maintaining the realism of original images. This functionality, essential for removing distracting objects in photos, has become a key feature in widely used applications like Photoshop, Google Photos, Microsoft Designer.
Equal contribution.  $\dagger$ Corresponding authors.  $\heartsuit$ Project leader.
@@
Figure 2: Comparison of the “mask-and-inpaint" paradigm and the proposed Masked-Region Guidance paradigm for object removal. Unlike the mask-and-inpaint approach, our paradigm uses the masked regions as critical guidance for object removal.
@@
Using the Masked-Region Guidance paradigm and the Syn4Removal dataset, we design a framework based on text-to-image diffusion models for object removal. To enhance the model’s robustness to varying mask shapes from user input, we introduce a mask enhancement technique that simulates different mask shapes over removal targets. Additionally, we incorporate CLIP-based visual guidance to assist the model in better understanding the removal targets. The resulting model called SmartEraser outperforms previous methods significantly in both quantitative and qualitative evaluations. We choose the name SmartEraser because the model can “smartly" identify removal targets and remove it while preserving its surrounding region.
@@ 3 SmartEraser
The goal of object removal is to completely erase the target object specified by a masked region and seamlessly generate consistent content based on the surrounding context. Most existing methods follow a “mask-and-inpaint" paradigm, where the masked area is excluded from the input, and the model attempts to inpaint the missing content based on the unmasked regions. While these methods can successfully remove objects in simple scenarios, they often regenerate new objects or produce blurs and artifacts in the masked region in more complex scenes.
In contrast, we propose a novel paradigm: Masked-Region Guidance. Instead of discarding masked regions, we retain it in the input, leveraging it to guide the removal process. This approach is based on the insight that, by knowing the exact removal targets, the model can precisely erase them while leaving the surrounding context maintained. In the following sections, we will first introduce the method of Masked-Region Guidance in Sec. 3.1. Then, we present Syn4Removal, a large-scale synthetic dataset designed for this paradigm in Sec. 3.2. Finally, we describe SmartEraser, a fine-tuned diffusion model based on the Mask-Region Guidance paradigm and Syn4Removal dataset in Sec. 3.3.
@@ 3.1 Masked-Region Guidance
Given an input image $\bm{x}$, and user-specified mask $\bm{m}$ containing the removal targets, the object removal task aims to produce the removal result $\bm{\hat{x}}$ while maintaining the realism of the original image. The mask $\bm{m}$ is a binary mask where 1 indicates the region containing the removal targets. Existing models [39, 35, 9, 55] are mostly based on the “mask-and-inpaint" paradigm, i.e., utilizing the unmasked region $\bm{x}\odot(\bm{1}-\bm{m})$ and mask $\bm{m}$ to predict the results without removal object in the masked region. These models are learned to maximize the following objective:
@@ 4.2 Comparasion with Previous Methods
ZITS++ [4] 53.440 0.838 0.820 0.388 0.522 15.442
MAT [24] 80.953 1.494 0.628 0.433 0.495 13.322
LaMa [39] 24.237 0.216 0.916 0.348 0.557 16.383
BLD [1] 73.086 1.760 0.834 0.471 0.467 15.775
RePaint [30] 71.542 1.639 0.812 0.381 0.510 15.861
SD-Inpaint [35] 69.502 0.324 0.857 0.369 0.537 16.111
CLIPAway [9] 25.458 0.123 0.915 0.333 0.577 17.434
PowerPaint [55] 24.058 0.294 0.926 0.308 0.602 18.100
SmartEraser 16.030 0.092 0.937 0.276 0.612 19.994
@@
ZITS++ [4] 8.153 0.229 0.899 0.350 0.634 19.453
MAT [24] 15.257 0.321 0.742 0.407 0.575 14.541
LaMa [39] 6.601 0.201 0.926 0.351 0.683 21.735
BLD [1] 24.407 0.746 0.913 0.418 0.632 19.346
RePaint [30] 12.314 0.822 0.924 0.325 0.644 20.607
SD-Inpaint [35] 11.119 0.196 0.912 0.320 0.656 21.691
CLIPAway [9] 8.320 0.193 0.925 0.323 0.666 21.953
PowerPaint [55] 6.582 0.275 0.934 0.278 0.715 23.870
SmartEraser 3.405 0.106 0.939 0.257 0.734 25.363
@@
ZITS++ [4] 5.932 0.193 0.810 0.358 0.581 17.315
MAT [24] 19.263 0.300 0.700 0.374 0.570 13.357
LaMa [39] 5.246 0.097 0.932 0.325 0.606 19.319
BLD [1] 20.286 0.705 0.896 0.464 0.525 16.539
RePaint [30] 12.677 0.696 0.908 0.378 0.553 17.689
SD-Inpaint [35] 11.528 0.107 0.897 0.348 0.572 18.105
CLIPAway [9] 5.530 0.093 0.930 0.344 0.593 18.561
PowerPaint [55] 5.025 0.200 0.931 0.291 0.663 20.594
SmartEraser 4.386 0.053 0.939 0.269 0.672 22.029
@@ 4.3 Ablation studies
Effect of the Masked-Region Guidance paradigm: We observe that the baseline model often regenerates objects instead of properly removing them, as it is trained in a traditional “inpainting" manner. Incorporating masked-region guidance significantly improves the removal performance by reducing the risk of object regeneration. However, due to the mismatch between the mask shape used during training and the user-defined mask during testing, the model does not always perfectly remove the target objects and preserve the surrounding context.